OpenStudy (anonymous):

what is a possible first step to eliminate the variable x in the following system? 2x-7y=3 4x+5y=-2

OpenStudy (anonymous):

multiply the first equation by -2 you get -4x+14y=-6 4x+5y=-2 cancel out the 4x's you get 19y=-8 19y/19=-8/19 y=-8/19

OpenStudy (lalaly):

never mind:D

OpenStudy (anonymous):

lol was about to say your suppose to add the two right...

OpenStudy (lalaly):

hehe yeh my bad:D

OpenStudy (anonymous):

it's all good.

OpenStudy (lalaly):

:D:D